OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem?

OpenStudy (anonymous):

total=112 g=x b=2x+16

OpenStudy (anonymous):

b = 16 + 2g g + 16 + 2g = 112 3g = 96 g = 32

OpenStudy (anonymous):

Nw find the no of boys

OpenStudy (anonymous):

b = 2g + 16

mathslover (mathslover):

put g = 32 and then solve for b

mathslover (mathslover):

b=2(32)+16 may i know what r u getting @Freedom17

OpenStudy (anonymous):

b = 2* 32 + 16

OpenStudy (anonymous):

just solve and find...

OpenStudy (anonymous):

3x+16=112 3x=112-16 x=96/3 =32

OpenStudy (anonymous):

@Freedom17 if u understand how to make equation then it may be easy

mathslover (mathslover):

let the user do it him/her self ..@best.shakir

OpenStudy (anonymous):

i had first given the hint

OpenStudy (anonymous):

@Freedom17 can u follow these all steps

OpenStudy (farmdawgnation):

@best.shakir Please don't provide solutions in the future after other users have already started working with the asker to try and engage them. We've already warned you a few times about this. You won't get another warning.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i did'nt gave the solution

OpenStudy (anonymous):

i had showed only the steps u check it

OpenStudy (anonymous):

@Yahoo! had written the ans already